OpenStudy (anonymous):

I need help with One-Variable Compound Inequalities. Please show your steps: -16 < less than or equal to 2+9x < greater than or equal to 11 Thank you!

OpenStudy (anonymous):

It is written this way it is -16 less than or equal to 2+9x < 11. The underline will not show up both parts is less than or equal to. -16 <2+9x < 11.

OpenStudy (johnweldon1993):

So in the long run...we want to solve for 'x' in the inequality... so we have \[-16 \le 2 + 9x \ge 11\] We want to get 9x by itself in the middle...so we need to subtract 2 from everything! why? because 2 + 9x = 2 = 9x....just 9x so... \[-16 -2 \le 2 + 9x - 2 \ge 11 - 2\] and this becomes \[-18 \le 9x \ge 9\] Now...we want to solve for 'x' right? not 9x...so lets divide EVERYTHING by 9...because 9x/9 = 1x...or just x right? so we have \[-18 /9 \le 9x / 9\ge 9/9\] \[? \le x \ge ?\] What are the 2 ?'s replaced with?

OpenStudy (johnweldon1993):

Wait...BOTH parts are less than or equal to???

OpenStudy (anonymous):

I got the first part -18 but did not get the 9’s. I thought everything was subtracted by 2 and divided by 2.

OpenStudy (anonymous):

4.5

OpenStudy (johnweldon1993):

Well then if THAT is correct....then in my entire explaination *it stays the same* just read it as \[\le x \le\] And no...everything IS subtracted by 2...*shown above* but then everything is divided by 9...after you subtract 2 from everything...it is gone! the only thing left is 9x

OpenStudy (anonymous):

9 is the denominator leaving 9x. If I graph this it is less than 9x?

OpenStudy (johnweldon1993):

Let me rewrite the steps...maybe that will clear it up

OpenStudy (anonymous):

Ok

OpenStudy (johnweldon1993):

\[-16 \le 2 + 9x \le 11\] \[-16 - 2 \le 2 - 2 + 9x \le 11-2\] \[-18 \le 9x \le 9\] With me so far? All I did was subtract EVERYTHING by 2 okay?

OpenStudy (anonymous):

Thanks John. Can you help with the second part of my equation? 12-x>15 or 4x-13>7

OpenStudy (johnweldon1993):

\[\frac{ -18 }{ 9 } \le \frac{ 9x }{ 9 } \le \frac{ 9 }{ 9 }\] This is me dividing EVERYTHING by 9...this will isolate 'x' after that..we'll have \[-2 \le x \le 1\] better?

OpenStudy (anonymous):

Thank you so very much! You are great!

OpenStudy (johnweldon1993):

Okay now onto that 2nd part....what do you think you would do first?

OpenStudy (anonymous):

I am thinking the opposition add instead of subtract?

OpenStudy (johnweldon1993):

12-x>15 or 4x-13>7 this means we have 2 equations to solve \[12 - x > 15\] \[4x-13 > 7\] Lets solve that first one to begin with.... what would you get?

OpenStudy (anonymous):

On the first answer when I draw the number line to reflect the answer what are my two points that I color in?

OpenStudy (johnweldon1993):

The first answer meaning the first problem we solved? *not the one we're working on now*?

OpenStudy (anonymous):

Yes I have to draw the answer on the number line also on both equations.

OpenStudy (johnweldon1993):

Okay...that very first problem we solved...we came up with \[-2 \le x \le 1\] So you can split this into 2 equations as well \[x \le 1\] and \[x \ge -2\] where would your 2 points be?

OpenStudy (anonymous):

Wow that is the answer I got but thought that it was incorrect! I am so happy you were still up to help me! Yeah!

OpenStudy (anonymous):

On the number line I would show -2 through 1, color it in at those points.

OpenStudy (johnweldon1993):

That would be correct...because if you say it outloud... 'x' is greater than or equal to -2 AND LESS than or equal to 1...so it would be all points between the 2

OpenStudy (anonymous):

Thank you that is what I wrote and I was just about to give up but thought I would ask for the correct steps and you provided that. Thank you so much! Can we solve the last two?

OpenStudy (johnweldon1993):

Absolutely....no like I wrote above...we have 2 equations \[12 - x > 15\] \[4x - 13 > 7\] What do you get when you solve that first one?

OpenStudy (anonymous):

Subtract 12 from all parts?

OpenStudy (johnweldon1993):

We can do that yes.... \[12 - 12 - x > 15 - 12\] \[- x > 3\] remember...we aren't solving for -x...we want 'x'...so what next?

OpenStudy (anonymous):

We do the same as above, subtract and then divide.

OpenStudy (johnweldon1993):

Right we subtracted 12 from both sides...*done above* and now we divide by '-' to solve for 'x' ****what do we do when we divide or multiply by a negative number in inequalities??****

OpenStudy (anonymous):

3

OpenStudy (anonymous):

I think when we multiply or divide we change the inequity to a positive number so we divide by 3, I think.

OpenStudy (johnweldon1993):

Not quite.... when we multiply or divide by a negative number....we flip the direction of the inequality sign...here I'll do this one for you \[-x > 3\] Divide both sides by -1 \[\frac{ -1x }{ -1 } > \frac{ 3 }{ -1 }\] this becomes \[x < -3\]

OpenStudy (anonymous):

This is a difficult concept but I am very happy to have your help.

OpenStudy (johnweldon1993):

No problem..I hope you understand it....we have 1 more equation to solve though remember? \[4x - 13 > 7\] what would you do first?

OpenStudy (anonymous):

I know I was just trying to figure out the one you just listed and my head is spinning. I do not understand so I am writing all of the steps down as you write them.

OpenStudy (johnweldon1993):

Want me to run over the steps again? I can probably explain it a little better! :)

OpenStudy (anonymous):

Yes, if you can walk me through the

OpenStudy (anonymous):

last one I will call it a day. the 4x<13<7

OpenStudy (anonymous):

I mean: 4x-13>7

OpenStudy (johnweldon1993):

Okay...and at the end I'll remind you about what I did for that 2nd one I just finished! so \[4x - 13 > 7\] what do we do?

OpenStudy (anonymous):

Subtract 4 from all parts and then divide.

OpenStudy (johnweldon1993):

Not quite...let's not touch the 'x' number....we want to solve for 'x' eventually so lets get rid of everything ELSE first

OpenStudy (anonymous):

Oh yeah we are going to flip the inequality symbol over so it points to the other direction.

OpenStudy (johnweldon1993):

Hang on hang on....lets take it step by step

OpenStudy (johnweldon1993):

\[4x - 13 > 7\] We don't touch the 'x' number here....we want to get rid of EVERYTHING else first... next to the 4x...we have a -13...how do we get rid of that?

OpenStudy (anonymous):

I cannot pretend to know the answer because I don't

OpenStudy (anonymous):

For the last two the assignment is asking for the solution set of both of the last equations with the number line shaded in the answer.

OpenStudy (johnweldon1993):

Okay...well just remember...since we want 4x to be alone....we want to get rid of the stuff next to it... \[4x - 13 > 7\] If we add 13 to both sides...what do we get? \[4x - 13 + 13 > 7 + 13\] notice how -13 + 13 = 0 right? so we'll have \[4x > 20\]

OpenStudy (anonymous):

so the answer is 4x is greater than 20 showing 4-20 on the number line for this one?

OpenStudy (johnweldon1993):

After that...we want to solve for x....not 4x....so we divide everything by 4 \[\frac{ 4x }{ 4 } > \frac{ 20 }{ 4 }\] this becomes \[x > 5\]

OpenStudy (anonymous):

x>5 is the final answer and the points to graph!

OpenStudy (johnweldon1993):

I really hope you're understanding this..... so we have *remember we started with 2 equations* our 2 answers are x < -3 and x > 5 so what points do you plot?

OpenStudy (anonymous):

The first set will show -3 on the graph the second will begin at 5 and go beyond that

OpenStudy (johnweldon1993):

Right....one part says "x is less than -3".....or "x is greater than 5" so the left hand would be an open circle at -3 and shading to the left....and the right side will have an open circle at 5 and shading to the right

OpenStudy (anonymous):

John, thank you so very much I have all ready tagged you to be a fan! I will write it all out for my assignment tomorrow. Thank you!

OpenStudy (johnweldon1993):

No problem! and if you need more or an explanation or anything...feel free to tag me in a post or message me!

OpenStudy (anonymous):

I will definitely do that. I don't know why it says that you are a novice, you seem like an expert!

OpenStudy (johnweldon1993):

Haha don't worry about that....I'm just so used to doing these problems from middle / high school...it is like second nature now...and one day soon...you will be able to do them like nothing as well :)

OpenStudy (anonymous):

LOL! I don't know about that nearly every assignment I turn in there is something wrong. To contact you in the future do I just click on your name above?

OpenStudy (johnweldon1993):

Yeah you can click on my name and click send message...or you could post a question and if you wanted just tag me by using the '@' button followed by my name

OpenStudy (anonymous):

Ok! I am so happy to have found you! LOL! Talk with you soon! Tuiti

OpenStudy (johnweldon1993):

Haha have a good night!

OpenStudy (anonymous):

You too!